OpenStudy (anonymous):

1.) Mrs. Computer is 3 times older than her daughter, Mousy. The sum of their ages is 52> How old is Mousy? 2.)Mike is 5 years more than twice as old as Tom. The sum of their ages is 65. How old is Mike?

OpenStudy (anonymous):

I hve a few more hang on lol

OpenStudy (anonymous):

1) given x+3x=52 means x=13 hence their ages are 13 and 39 respectively

OpenStudy (anonymous):

2) here x+(2x+5) =65 or 3x=60 or x=20 their ages are 20 and 45 respectively

OpenStudy (anonymous):

Find three consecutive integers such that the sum of the second and the third is 17. Which equation would be used to solve this word problem? a.)x + (x + 1) + (x + 2) = 17 b.)(x + 1) + (x + 2) = 17 c.)x + (x + 2) + (x + 4) = 17 d.)2(x + 2) + 3(x + 4) = 17 The sum of three consecutive integers is 108. Which equation would be used to solve this word problem? (HINT: The equation has been simplified some.) a.)3x + 3 = 108 b.)3x = 108 c.)3(x+3) = 108 d.)3x + 1 = 108 Find two consecutive odd integers whose sum is 72. Which equation would be used to solve this problem? (HINT: The equation has been simplified some.) a.)2x = 72 b.)2x + 2 = 72 c.)x + 2 = 72 d.)2x + 1 = 72

OpenStudy (anonymous):

17 b 18a 19 b

OpenStudy (anonymous):

thankyou so much! word problems always get me

OpenStudy (anonymous):

welcome